OpenStudy (anonymous):

Solve for x, given the equation Square root of x minus 5 + 7 = 11. x = 21, solution is extraneous x = 21, solution is not extraneous x = 81, solution is extraneous x = 81, solution is not extraneous

OpenStudy (anonymous):

@kawaiicat123

OpenStudy (anonymous):

x = 21, solution is extraneous x = 21, solution is not extraneous x = 81, solution is extraneous x = 81, solution is not extraneous

OpenStudy (anonymous):

?????

OpenStudy (raffle_snaffle):

x-5+7=11

OpenStudy (anonymous):

Sorry. Wrong thing.

OpenStudy (raffle_snaffle):

5+7=?

OpenStudy (anonymous):

12

OpenStudy (anonymous):

???? @kawaiicat123 please help me

OpenStudy (anonymous):

I'll try.

OpenStudy (anonymous):

Subtract 7 from both sides and then stop to think about it.

OpenStudy (anonymous):

Subtract 7 from each side. 7 - 7 = ?? 11 - 7 = ??

OpenStudy (anonymous):

0 and 4